# Perception-aware Planning for Quadrotor Flight in Unknown and Feature-limited Environments

### v1

@@ Abstract
Various studies on perception-aware planning have been proposed to enhance the state estimation accuracy of quadrotors in visually degraded environments. However, many existing methods heavily rely on prior environmental knowledge and face significant limitations in previously unknown environments with sparse localization features, which greatly limits their practical application. In this paper, we present a perception-aware planning method for quadrotor flight in unknown and feature-limited environments that properly allocates perception resources among environmental information during navigation. We introduce a viewpoint transition graph that allows for the adaptive selection of local target viewpoints, which guide the quadrotor to efficiently navigate to the goal while maintaining sufficient localizability and without being trapped in feature-limited regions. During the local planning, a novel yaw trajectory generation method that simultaneously considers exploration capability and localizability is presented. It constructs a localizable corridor via feature co-visibility evaluation to ensure localization robustness in a computationally efficient way. Through validations conducted in both simulation and real-world experiments, we demonstrate the feasibility and real-time performance of the proposed method. The source code will be released to benefit the community
@@ I Introduction
Visual-inertial odometry (VIO) has been a popular method for state estimation of unmanned aerial vehicles (UAVs) due to its lightweight and low-power characteristics. However, even well-established methods [1, 2], struggle in feature-scarce conditions such as textureless surfaces and low light. To mitigate localization failures, researchers have carried out numerous studies of perception-aware planning, incorporating perception constraints into trajectory generation to enhance localizability.
@@ III Problem Formulation
During the navigation, the constraints of safety, dynamic feasibility and localizability must be strictly satisfied. Specifically, the following situations will be judged as mission failures:
@@ V-A1 Viewpoint Transition Graph
where ${L}$ represents all currently known features, $\mathbb{I}(\cdot)$ is an indicator function that equals 1 if the condition is met and 0 otherwise, and $\text{FOV}(\cdot)$ is the field of view function, which returns the set of points visible from a given viewpoint. The thresholds $\mathcal{V}_{thr}$ and ${FC}_{thr}$ denote the minimum number of features and frontier cells that must be visible for any viewpoint sampled. An example of frontier-based viewpoints is shown in Fig.3(b).
Feature-based Viewpoints: Feature-based viewpoints are also considered into the graph, which are generated based on the following idea: all currently known features ${L}$ can be clustered into several groups, resulting in feature clusters ${LC}_{1},{LC}_{2},\dots,{LC}_{N}$, with each cluster containing enough features for localization. A feature-based viewpoint is expected to observe multiple feature clusters, such as ${LC}_{a}$ and ${LC}_{b}$. In this case, there ensures a transition from the viewpoint observing $LC_{a}$ to the one observing $LC_{b}$ through the feature-based viewpoint, while maintaining localizability in the path.
@@
where $v_{\max}$ and $\dot{\psi}_{max}$ denote the maximum translational and yaw velocities, respectively.
@@ V-C2 Trajectory Optimization
  • $\mathbf{c}_{j}$ remains invisible to $\mathbf{s}(\psi_{i}^{{}^{\prime}}),\forall\psi_{i}^{{}^{\prime}}\in(\mathfrak{l}_{i},\mathfrak{u}_{i})$
@@ VII Conclusion
This paper presents a perception-aware planning method for quadrotor flight in unknown and feature-limited environments. Our approach establishes a systematic mechanism to properly allocate perception resources among environmental information in a computationally efficient way. Firstly, a viewpoint transition graph is proposed to adaptively select local targets, balancing localization and navigation efficiency. Besides, at the trajectory generation level, a localizable corridor-based method is proposed to ensure the localization robustness efficiently without numerous evaluation of visibility during the optimization. Our approach enables efficient perception-aware navigation without any prior environmental knowledge. Both simulations and real-world experiments are conducted, and the results verify the feasibility of our method.

### v2

@@ Abstract
Various studies on perception-aware planning have been proposed to enhance the state estimation accuracy of quadrotors in visually degraded environments. However, many existing methods heavily rely on prior environmental knowledge and face significant limitations in previously unknown environments with sparse localization features, which greatly limits their practical application. In this paper, we present a perception-aware planning method for quadrotor flight in unknown and feature-limited environments that properly allocates perception resources among environmental information during navigation. We introduce a viewpoint transition graph that allows for the adaptive selection of local target viewpoints, which guide the quadrotor to efficiently navigate to the goal while maintaining sufficient localizability and without being trapped in feature-limited regions. During the local planning, a novel yaw trajectory generation method that simultaneously considers exploration capability and localizability is presented. It constructs a localizable corridor via feature co-visibility evaluation to ensure localization robustness in a computationally efficient way. Through validations conducted in both simulation and real-world experiments, we demonstrate the feasibility and real-time performance of the proposed method. The source code is released for the reference of the community [^1] https://github.com/Robotics-STAR-Lab/LA-Planner.
@@ I Introduction
Visual-inertial odometry (VIO) has been a popular method for state estimation of unmanned aerial vehicles (UAVs). However, even well-established methods [1, 2], struggle in feature-scarce conditions such as textureless surfaces and low light. To mitigate localization failures, researchers have carried out numerous studies of perception-aware planning, incorporating perception constraints into trajectory generation to enhance localizability.
@@ III Problem Formulation
During the navigation, the constraints of safety, dynamic feasibility, and localizability must be strictly satisfied. Specifically, the following situations will be judged as failures:
@@ V-A1 Viewpoint Transition Graph
[…]
Feature-based Viewpoints: Feature-based viewpoints are also considered into the graph, which are generated based on the following idea: all currently known features ${L}$ can be clustered into several groups, resulting in feature clusters ${LC}_{1},{LC}_{2},\dots,{LC}_{N}$, with each cluster containing enough features for localization. A feature-based viewpoint is expected to observe multiple feature clusters, such as ${LC}_{i}$ and ${LC}_{j}$. In this case, a transition from the viewpoint observing ${LC}_{i}$ to that observing ${LC}_{j}$ is ensured via the feature-based viewpoint, while maintaining localizability along the path.
@@
where $v_{\max}$ and $\dot{\psi}_{max}$ denote the maximum translational and yaw velocities, respectively. The cost is the minimum time for both motions in parallel, set by the slower one.
@@ VII Conclusion
This paper presents a perception-aware planning method for quadrotor flight in unknown and feature-limited environments. Our approach establishes a systematic mechanism to properly allocate perception resources among environmental information in a computationally efficient way. Firstly, a viewpoint transition graph is proposed to adaptively select local targets, balancing localization and navigation efficiency. Besides, at the trajectory generation level, a localizable corridor-based method is proposed to ensure the localization robustness efficiently without numerous evaluation of visibility during the optimization. Our approach enables efficient perception-aware navigation without any prior environmental knowledge. The feasibility of our method has been validated through both simulations and real-world experiments.